# Stochastic Sampling of Operator Growth Dynamics

### v1

@@ Abstract
We put forward a Monte Carlo algorithm which samples the Euclidean time operator growth dynamics at infinite temperature. Crucially, our approach is free of the numerical sign problem for a range of quantum many-body spin systems, allowing for numerically exact and unbiased calculations. We apply this methodological headway to study the high-frequency dynamics of the mixed-field quantum Ising model (QIM) at one and two dimensions. Our results corroborate the recently proposed operator growth hypothesis, predicting an exponential fall-off of generic response functions at large frequencies. Remarkably, our calculations are sufficiently sensitive to detect subtle logarithmic corrections of the hypothesis in one dimension. In addition, in two dimensions, we uncover a non-trivial crossover between two large frequency decay rates. Lastly, we find that the growth rate of the operator support size is dominated by boundary effects, with a logarithmic correction in one dimension.
Introduction — Within the Heisenberg representation of quantum mechanics, the time evolution of operators is governed by their commutation relations with the Hamiltonian. Under generic many-body dynamics, local operators are expected to evolve into increasingly complex and non-local operator strings. This process of “operator growth” is equivalent to the scrambling of quantum information in complex systems, and closely related to quantum chaos. As such, understanding the precise properties of operator growth is a fundamental line of research that has driven a great amount of theoretical and experimental work [1, 2, 3, 4, 5, 6, 7, 8, 9]. It has direct links to the question of how the laws of quantum mechanics bring about thermalization or its failure in isolated systems [10, 11, 12, 13, 14]. It is also a useful tool for investigating the fate of information in a black hole and how to retrieve it [15, 16, 17, 18, 19].
In this context, an interesting way to quantify the rate of operator growth, besides the much studied out-of-time order correlator, is to examine the high-frequency regime of the spectral density associated with generic correlators [20]. Recently, Ref. [21] proposed an operator growth hypothesis that translates to an exponentially decaying large frequency tail
@@
Testing the operator growth hypothesis in concrete lattice models is particularly challenging due to the inherent difficulty of simulating many-body quantum dynamics with classical resources. Direct evidence for the hypothesis is largely limited to large-N or semi-classical models [23, 30, 33, 20, 34] and brute-force numerics of small to intermediate system sizes that suffer from sizable finite size effects [35]. Although analytical bounds supporting the conjecture have been obtained for finite-dimensional systems [36, 37], these results are not satisfactory in that they do not provide an accurate estimate of $\omega_{0}$ or the precise functional form of the spectral function. It is, therefore, desirable to devise efficient numerical schemes that can reliably test the operator growth hypothesis.
Figure 1: (a) A space-time snapshot of the Euclidean time operator growth process extracted from the $400^{\text{th}}$ moment of the Pauli $Z$ operator at site 0 of the 1D mixed field QIM. Transformations are generated by Liouvillians between time slices (indexed by $k$). The colors correspond to the different Pauli operators. (b) The average occupation probability on a site by a non-identity Pauli operator resulting from operator growth arising from the same Monte Carlo dynamics as (a).
@@
In this Letter, we present a Monte Carlo method that samples the Euclidean time operator growth dynamics at infinite temperature in an unbiased and statistically exact manner. The Monte Carlo method tracks the Euclidean time evolution of operators (Fig. 1a) instead of the standard wave function approach that is common in conventional quantum Monte Carlo calculations. Crucially, our scheme does not suffer from the notorious numerical sign problem when applied to a broad family of lattice spin models of interest. We employ our method to study the mixed-field QIM in both one and two dimensions. Our results provide high-precision tests of the hypothesis. In particular, we pinpoint the subtle log-correction to Eq. 1 in one dimension, and unveil an intriguing crossover between two high-frequency tails in two dimensions. Finally, we determine the dynamical scaling properties of the operator support size.
@@
Operator-space partition function — In statistical physics, Monte Carlo schemes aim to sample from a partition function representing a weighted sum over the configuration space. In our case, the relevant partition function is the auto-correlation function of a local operator, $O$, analytically continued to imaginary time:
$\mathcal{Z}(\tau):=(O|e^{\mathcal{L}\tau}O)\,.$ (2)
Here, $\mathcal{L}O=[H,O]$ is the Liouvillian super operator representing the commutation with the Hamiltonian $H$, and $(A|B)=\mathrm{Tr}[A^{\dagger}B]/\mathrm{Tr}[1]$ denotes the infinite-temperature scalar product on operator space. We shall also assume that $O$ equals a product of one-site operators. To make contact with the real frequency spectral density (and especially its large frequency tails), it is useful to consider the Taylor expansion of $Z(\tau)$ at $\tau=0$:
$\mathcal{Z}(\tau)=\sum_{n=0}^{\infty}\mu_{2n}\frac{\tau^{2n}}{(2n)!}$ (3)
@@
In the last equation, we identify the Taylor expansion coefficients $\{\mu_{2n}\}$ with the moments of the spectral function $\Phi(\omega)$, which follows from a Fourier transform of the analytic continuation to real-time correlators, $\mathcal{C}(t)=\mathcal{Z}(\tau=it)$. Thus, high moments encode information about the high-frequency tail of $\Phi(\omega)$. In particular, an exponential tail (Eq. 1) yields:
@@
for large $n$. We recall that the moments grow rapidly, such that $\mathcal{Z}(\tau)$ converges only when $\tau<\tau_{*}=1/\omega_{0}$, and has a pole at $\tau_{*}$.
To evaluate the moments via stochastic sampling, we express $\mu_{2n}$ as a sum over paths in operator space. Assuming local dynamics, we write $\mathcal{L}=\sum_{i,a}\mathcal{L}_{i,a}$ , with $\left\{\mathcal{L}_{i,a}\right\}$ denoting local Liouvillian of type $a$ acting on site $i$ , as dictated by the underlying Hamiltonian, $H=\sum_{i,a}H_{i,a}$, with $\mathcal{L}_{i,a}O=[H_{i,a},O]$. Following the above definitions, we expand the product $(O|\mathcal{L}^{2n}|O)$, by introducing operator space resolution of identities between individual Liouvillians:
@@
where the sum is over all local Liouvillian sequences, specified by $2n$ position-type pairs, $\mathcal{S}_{2n}=\{(i_{k},a_{k})\}_{k=0}^{2n-1}$, with nonzero weight $W_{\mathcal{S}_{2n}}\neq 0$. This construction implicitly defines operator histories $(O_{0}=O,O_{1},\dots,O_{2n}=O_{0})$ and is reminiscent of the Stochastic Series Expansion (SSE) [38, 39]. We emphasize that, unlike standard quantum Monte Carlo approaches, here, the configuration space is defined via the space of local operators, and transitions between them are generated by the Liouvillian.
The weights $W_{\mathcal{S}}$ are generally non-positive and can lead to a sign problem, even in cases where the associated Hamiltonian problem is sign problem free. Remarkably, for a class of nontrivial models, the sum in Eq. 6 contains strictly non-negative real terms, $W_{\mathcal{S}}\geq 0$ [37]. This class includes the mixed-field QIM, with the Hamiltonian,
@@
Here, $\left\langle{i,j}\right\rangle$ define nearest neighbours on a hypercubic lattice, and $X_{i},Z_{i}$ are the Pauli operators act on sites $i$ along the $x,z$ spin axes, respectively. In what follows, we will focus on the numerical study of the mixed field QIM with positive and uniform $J$, $h^{Z}$, $h^{X}$.
Figure 2: A typical configuration contributing to the $4$-th moment with the horizontal and vertical axes being temporal and spatial respectively. The sites are populated by the $X$, $Y$, and $Z$ Pauli operators or the $I$ (identity) operator. The temporal direction is generated by the Liouvillians of type $\mathcal{L}_{Z}$, $\mathcal{L}_{ZZ}$ and $\mathcal{L}_{X}$ corresponding to the three types of local Hamiltonian terms. The empty Liouvillian slots correspond to null superoperators.
Numerical methods and observables — Following the expansion of $\mathcal{Z}(\tau)$ in Eq. 3, we now turn to devise a Monte Carlo algorithm to numerically estimate the moments $\mu_{2n}$. To that end, we set $\tau=1/J$ in Eq. 2 and restrict the expansion order to obey $2n\leq 2n_{\max}$. Analogously to SSE, we consider Liouvillian sequences, $\mathcal{S}_{2n_{\max}}$, of fixed length $2n_{\max}$, which for an expansion order $n$ include additional $(2n_{\max}-2n)$ factors of the null super-operator $\mathcal{L}_{0,0}\equiv I$.
We note that the above truncation is natural in the context of imaginary time Hamiltonian dynamics via SSE simulations. There, for given values of microscopic parameters and temperature, the probability distribution of the expansion order rapidly decays away from its mean value. By contrast, here, we will be interested in targeting all moments $\mu_{2n}$ for $2n\leq 2n_{\max}$. As will become clearer below, this entails visiting each expansion order at roughly equal frequency.
The resulting partition function has the following form,
$G=$ $\sum_{\mathcal{S}_{2n_{\max}}}\frac{1}{J^{2n}(2n)!}$ (8)
×$\times\frac{(2n_{\max}-2n)!(2n)!}{(2n_{\max})!}\prod_{k=0}^{2n_{\max}-1}(O_{k+1}|\mathcal{L}_{i_{k},a_{k}}|O_{k}),$
where the combinatorial factors arise from the redundancy in the assignment of null operators. Eq. 8 has a natural diagrammatic interpretation as a sum over $D=d+1$ dimensional space-time configurations consisting of on-site Pauli operator, with the dynamics generated by insertions of local Liouvillian super-operators, transitioning between Pauli strings, see Fig. 2.
@@
In the following, we present a set of local Monte Carlo moves that enable traversing between all allowed configurations: (1) add-drop – allows switching between different $n$ sectors by adding or removing a pair of identical local Liouvillians located on the same sites at two consecutive time slices. (2) swap – swaps two commuting Liouvillians at different time slices, without changing the expansion order $n$. This move is proposed only if there are no non-commuting Liouvillians at intermediate time slices. (3) interact – here, two temporally consecutive $\mathcal{L}_{ZZ}$ Liouvillians that share a single spatial site “interact” to create a pair of temporally consecutive $\mathcal{L}_{Z}$ Liouvillians acting on next to nearest neighbour sites. This move allows accessing configurations with an odd number of $\mathcal{L}_{ZZ}$ and $\mathcal{L}_{Z}$ Liouvillians acting on a given site.
In practice, we found it essential to apply a reweighting approach, in the spirit of the Wang-Landau algorithm [40, 41, 42], in order to ensure ergodicity due to rapid fall off (or rise) of the expansion order distribution towards large moments. This is achieved by iteratively tuning the relative visiting rate of each expansion order until reaching an approximately uniform distribution across expansion order sectors. In other words, we sample the partition sum $\mathcal{Z}(\tau)$ near criticality, $\tau\nearrow\tau_{*}$. Additional technical details on the Monte Carlo algorithm are relegated to [43].
@@
$r_{n}^{2}:=\frac{\mu_{2n+2}}{\mu_{2n}}=\frac{\left\langle\delta_{n^{\prime},n+1}\right\rangle_{n^{\prime}}}{\left\langle\delta_{n^{\prime},n}\right\rangle_{n^{\prime}}}\times J^{2}(2n+2)(2n+1).$ (9)
In the above equation, averaging is carried out over the probability distribution of the expansion order $P(n^{\prime})$, where the Kronecker delta $\delta_{n^{\prime},n}$ counts Monte Carlo events satisfying $n^{\prime}=n$. Effectively, $r^{2}_{n}$ is tied to the ratio between the visiting frequencies of expansion orders $n+1$ and $n$. The large frequency behavior of Eq. 5, predicts a linear growth of $r_{n}$ with $n$,
$r_{n}\simeq\frac{2\omega_{0}n}{e}.$ (10)
In addition, the prefactor allows a numerical estimation of the decay rate $\omega_{0}$.
@@
Figure 3: (a) $r_{n}$ vs $n$ for the $Z$ operator with $J=h^{Z}=h^{X}=1$ in 1D. The inset zooms in on the first 35 moments, comparing our calculations with exact brute force numerics. (b) $\Delta r_{n}$ vs $n$ for 1D. This asymptotically falls to zero. Multiplying by $\log(n)$ recovers a constant plateau. (c) Same as (a) for 2D. The inset has the first 16 moments compared against exact computation. (d) $\Delta r_{n}$ vs $n$ for 2D. Unlike 1D here the $\log$ multiplication isn’t needed. The inset shows $n\Delta r_{n}$ vs $n$.
Numerical Results – In the following, we present Monte Carlo data for the $Z$ operator dynamics in the mixed field QIM with the microscopic parameters $h^{X}=h^{Z}=J=1$ in 1D and 2D. We begin our presentation by examining the 1D case. We recall that, due to the restricted dynamics in 1D, the linear growth in Eq. 10 receives a logarithmic correction such that $r_{n}\sim n/\log{n}$ ) [44, 36, 32, 45, 21]. In Fig. 3a, we depict $r_{n}$ as a function of $n$. We first note the excellent agreement with brute force calculation (suffering from an unfavourable exponential scaling with the expansion order) as shown in the inset, serving as a non-trivial benchmark of our approach.
Furthermore, we observe that the large $n$ growth of $r_{n}$ deviates from a linear trend. To pin down the source of this deviation and check whether it can be attributed to the aforementioned logarithmic correction in Fig. 3b, we plot $\Delta r_{n}\times\log(n)$, where $\Delta r_{n}$ is a numerical derivative extracted using the five-point stencil method. This quantity is expected to reach a plateau at large $n$ with the logarithmic correction. Indeed, we find that only in the presence of the log multiplication, the desired plateau is obtained. By contrast, $\Delta r_{n}$ asymptotically decays to zero due to the $1/\log(n)$ factor. We note that this is the first numerically exact confirmation for the predicted log correction associated with the restricted 1D dynamics, as previous brute-force numerics cannot resolve this subtle functional dependence convincingly due to the limited number of accessible moments [21, 35, 46].

### v2

@@ Abstract
We put forward a Monte Carlo algorithm that samples the Euclidean time operator growth dynamics at infinite temperature. Crucially, our approach is free from the numerical sign problem for a broad family of quantum many-body spin systems, allowing for numerically exact and unbiased calculations. We apply this methodological headway to study the high-frequency dynamics of the mixed-field quantum Ising model (QIM) in one and two dimensions. The resulting quantum dynamics display rapid thermalization, supporting the recently proposed operator growth hypothesis. Physically, our findings correspond to an exponential fall-off of generic response functions of local correlators at large frequencies. Remarkably, our calculations are sufficiently sensitive to detect subtle logarithmic corrections of the hypothesis in one dimension. In addition, in two dimensions, we uncover a non-trivial dynamical crossover between two large frequency decay rates. Lastly, we reveal spatio-temporal scaling laws associated with operator growth, which are found to be strongly affected by boundary contributions.
I Introduction
Within the Heisenberg representation of quantum mechanics, the time evolution of operators is governed by their commutation relations with the Hamiltonian. Under generic many-body dynamics, local operators are expected to evolve into increasingly complex and non-local operator strings. This process of “operator growth” is equivalent to the scrambling of quantum information in complex systems and is closely related to quantum chaos. As such, understanding the precise properties of operator growth is a fundamental line of research that has driven a great amount of theoretical and experimental work [1, 2, 3, 4, 5, 6, 7, 8, 9]. It has direct links to the question of how the laws of quantum mechanics bring about thermalization or its failure in isolated systems [10, 11, 12, 13, 14]. It is also a useful tool for investigating the fate of information in a black hole and how to retrieve it [15, 16, 17, 18, 19].
In this context, an interesting way to quantify the rate of operator growth, besides the much studied out-of-time order correlator, is to examine the high-frequency regime of the spectral density $\Phi(\omega)$ associated with generic dynamical correlators [20]. Recently, Ref. [21] proposed an operator growth hypothesis that translates to an exponentially decaying large frequency tail
@@ I Introduction
Figure 1: (a) The operator growth history of the $Z$ Pauli operator associated with a typical sequence belonging to $\mathcal{S}_{2n}$ for $n=200$ in Eq. 6 in the Pauli basis, for the mixed field QIM with Hamiltonian given in Eq. 9. The horizontal axis corresponds to space, and the vertical axis is generated by the temporal index $k$. The colors correspond to different Pauli operators at each site. (b) The average Pauli operator density across growth histories contributing to $\mu_{2n}$ for $n=200$. Here, we only distinguish between the non-trivial Pauli operators and the identity. Both (a) and (b) have been extracted from our Monte Carlo simulations.
@@
Studying operator growth dynamics in concrete lattice models is particularly challenging due to the inherent difficulty of simulating many-body quantum dynamics with classical resources, as reflected by the dynamical sign problem and the unbounded growth of entanglement entropy in thermalizing systems [33, 34]. From the perspective of the operator growth hypothesis, direct evidence is largely limited to large-N or semi-classical models [23, 30, 35, 20, 36] and brute-force numerics of small to intermediate system sizes that suffer from sizable finite size effects [37]. Although analytical bounds supporting the conjecture have been obtained for finite-dimensional systems [38, 39], these results are not satisfactory in that they do not provide an accurate estimate of $\omega_{0}$ or the precise functional form of the spectral function. It is, therefore, desirable to devise efficient numerical schemes that can reliably test the operator growth hypothesis.
@@
In this article, we present a Monte Carlo method that samples the Euclidean time operator growth dynamics at infinite temperature in an unbiased and statistically exact manner. A unique feature of our formalism is that it tracks the Euclidean time evolution of operators (Fig. 1a) instead of the standard wave function approach that is common in conventional Quantum Monte Carlo (QMC) calculations. Crucially, our scheme does not suffer from the notorious numerical sign problem when applied to a broad family of lattice spin models of interest. By way of example, we employ our method to study the mixed-field Quantum Ising Model (QIM) in one- and two-dimensions and succeed in computing an order of magnitude more number of spectral function moments than what is accessible using brute force methods. This progress gives us direct access to high-frequency decay rates of the spectral tail. Furthermore, our results provide high-precision tests of the operator growth hypothesis. In particular, we pinpoint the subtle log-correction to Eq. 1 in 1D and unveil non-universal spectral features like an intriguing crossover between two high-frequency tails in 2D. Additionally, we determine the dynamical scaling properties of the operator support size.
II Constructing the Operator-Space Partition Function
@@
In statistical physics, Monte Carlo schemes aim to tackle exponentially hard computational problems by mapping them to a sampling problem from a probability distribution. A useful starting point for constructing our partition function is to examine the auto-correlation function of a local operator, $O$, analytically continued to imaginary time $t=i\tau$:
$\mathcal{C}(\tau)=\langle O(\tau)O(0)\rangle=(O|e^{\mathcal{L}\tau}|O).$ (2)
[…]
To make contact with the real frequency spectral density (and especially its large frequency tails) we Taylor expand $\mathcal{C}(\tau)$ about $\tau=0$:
$\mathcal{C}(\tau)=\sum_{n=0}^{\infty}\mu_{2n}\frac{\tau^{2n}}{(2n)!}$ (3)
@@
[…]
@@
for large $n$. The moments grow rapidly, such that $\mathcal{C}(\tau)$ converges only when $\tau<\tau_{*}=1/\omega_{0}$, and has a pole at $\tau_{*}$.
We aim to extract dynamical information by evaluating the moments. To map this calculation to a statistical mechanics problem, we express $\mu_{2n}$ as a sum over paths in operator space. Assuming local dynamics, we write $\mathcal{L}=\sum_{i,a}\mathcal{L}_{i,a}$ , with $\left\{\mathcal{L}_{i,a}\right\}$ denoting local Liouvillian of type $a$ acting on site $i$, as dictated by the underlying Hamiltonian, $H=\sum_{i,a}H_{i,a}$, with $\mathcal{L}_{i,a}|O)=|[H_{i,a},O])$. We then expand $(O|\mathcal{L}^{2n}|O)$ by introducing operator space resolution of identities between individual Liouvillians:
@@
where the sum is over all local Liouvillian sequences, specified by $2n$ position-type pairs, $\mathcal{S}_{2n}=\{(i_{k},a_{k})\}_{k=0}^{2n-1}$, with nonzero weight $W_{\mathcal{S}_{2n}}\neq 0$. This construction implicitly defines operator histories between orthogonal operators $(O_{0}=O,O_{1},\dots,O_{2n}=O_{0})$. Eq. 6 imparts moments $\mu_{2n}$ the interpretation of a sum over all $2n$ long operator string paths generated by local Liouvillian sequences $S_{2n}$ such that $O_{0}=O_{2n}=O$. Fig. 1 illustrates this interpretation. We now consider the sum over all such operator paths up to $2n_{\max}$ in length:
$G=\sum_{n=0}^{n_{\max}}\mu_{2n}=\sum_{n=0}^{n_{\max}}\sum_{\mathcal{S}_{2n}}\prod_{k=0}^{2n-1}(O_{k+1}|\mathcal{L}_{i_{k},a_{k}}O_{k}).$ (7)
One can view $G$ as a partition function associated with a probability distribution on $n$ with $P(n)=\mu_{2n}/G$. In that case, the relative frequencies of sampling different values of $n$ from $P(n)$ correspond to the relative magnitudes of the moments. This forms the central idea behind our sampling approach to computing the moments.
To tailor $G$ for numerical Monte Carlo calculations, we make additional adjustments. Firstly, we make each operator string path $2n_{\max}$ long by inserting additional factors of identity Liouvillians. This results in duplication of paths which is compensated for by dividing by a binomial combinatorial factor. This step is reminiscent of the Stochastic Series Expansion (SSE) [40, 41], except that here, the configurations are in operator space with transitions being generated by local Liouvillians as opposed to local Hamiltonians. Also in contrast to SSE, where the probability distribution of the expansion order rapidly decays away from its mean value (associated with the energy), here we will be interested in sampling all moments $\mu_{2n}$ and thus drawing samples from all orders of the expansion. Eq. 5 however suggests that the moments grow rapidly with $n$. Naively sampling from $G$ would thus result in a highly skewed sampled histogram, with almost all samples coming from the large expansion orders close to $n_{\max}$. This would result in poor statistics for small to intermediate expansion orders. To counter this, we reweigh the moments by numerically determined factors $w_{n}$ such that $w_{n}\mu_{2n}$ is of the same order of magnitude for all $n$. Accounting for the reweighting and the insertion of identity Liouvillians, our final generalized partition function now takes the form:
$\mathcal{Z}=\sum_{n=0}^{n_{\max}}w_{n}\underbrace{\sum_{\mathcal{S}_{2n}^{\prime}}\frac{1}{{2n_{\max}\choose 2n}}\prod_{k=0}^{2n_{\max}-1}(O_{k+1}|\mathcal{L}_{i_{k},a_{k}}O_{k})}_{\mu_{2n}}.$ (8)
Here, now, $\mathcal{S}_{2n}^{\prime}$ sums over all possible $2n_{\max}$ long Liouvillian sequences with $2n-2n_{\max}$ Liouvillian identities and (${2n_{\max}\choose 2n}$ is the aforementioned combinatorial binomial factor, compensating for the over counting introduced by identity Liouvillians.
It should be noted that in $n_{\max}$ steps of commutation with a $k$ local Hamiltonian, an operator can spatially grow by at most $(k-1)\times n_{\max}$ sites. Thus, for an operator $O$ with finite support, a lattice with a linear spatial size larger than the sum of the support size of $O$ and $(k-1)\times n_{\max}$ avoids any finite size effects in the calculation of the first $n_{\max}$ moments.
An interesting observation is that for a particular choice of weights $w_{n}=\tau^{2n}/(2n)!$, $\mathcal{Z}$ becomes the truncated Taylor expansion of $\mathcal{C}(\tau)$ from Eq. 3. For large $n_{\max}$, a choice of weights $w_{n}$ that ensures a flat distribution across $n$ would correspond to $\mathcal{C}(\tau)$ near criticality, $\tau\nearrow\tau_{*}$.
Figure 2: A typical configuration contributing to $\mathcal{Z}$ in Eq. 8 with $n=4$, $n_{\max}=5$ and $L=4$ for the 1D mixed field QIM with Hamiltonian in Eq. 9. The horizontal and vertical axes correspond to the temporal index $k$ and the spatial index $i_{k}$ respectively. The sites are populated by the $X$, $Y$, and $Z$ Pauli operators or the $I$ (identity) operator. The temporal evolution is generated by Liouvillians of type $\mathcal{L}_{Z}$, $\mathcal{L}_{ZZ}$ and $\mathcal{L}_{X}$ arising from the longitudinal field, bond and transverse field terms in the Hamiltonian respectively. The empty Liouvillian slots correspond to identity Liouvillians.
The weights $W_{\mathcal{S}}$ in Eq. 6 arise out of a product of commutations of operators and are thus generally non-positive even when the associated Hamiltonian problem is sign-problem free. Remarkably, for a class of nontrivial models, the sum in Eq. 6 contains strictly non-negative real terms, $W_{\mathcal{S}}\geq 0$ [39], see further details in Section III. This class includes the mixed-field QIM, with the Hamiltonian:
@@
Here, $\left\langle{i,j}\right\rangle$ define nearest neighbours on a hypercubic lattice and $X_{i},Z_{i}$ are the Pauli operators acting on sites $i$ along the $x,z$ spin axes, respectively. For such spin $1/2$ models, Pauli strings, i.e. operators formed by the tensor product of local Pauli operators, are a convenient orthogonal basis choice for the resolution of Identities in Eq. 6. To avoid the sign problem, it is sufficient to assume that operator $O_{0}$ is itself a Pauli string. This then imparts the configurations in Eq. 8 a natural diagrammatic interpretation as a sum over configurations on a $D=d+1$ dimensional space-time lattice with dimensions $L^{d}\times(2n_{\max}+1)$. The lattice consists of Pauli strings spanning the spatial dimensions with the temporal dimension generated by local Liouvillians connecting Pauli strings. See Fig. 2.
III Sign Problem Resolution
Figure 3: The operator state graph of a site in the mixed field QIM Hamiltonian. The different states correspond to Pauli matrices, transitions between which are generated by local Liouvillians. The red and green arrows depict transitions arising from commutators with $+i$ and $-i$ phases respectively. The black arrows indicate transitions arising from matrix multiplications as can arise from $\mathcal{L}_{ZZ}$ commutators and thus have no associated phase.
To understand the resolution of the sign problem, it is useful to consider the operator state graph for a single site on the Pauli string lattice. Fig. 3 illustrates the same for the mixed field QIM. A site starts off in the state corresponding to the initial Pauli string $O_{0}$. As the Pauli string evolves, the site traces a path in the operator state graph. The boundary conditions imposed in Eq. 6 however demand that each site return to the same state it began in. It can be seen from Fig. 3 that in the mixed field QIM, for a site to return to the same state, it must encounter the same number of red transitions as green transitions, each associated with a phase factor of $+i$ and $-i$, respectively. Consequently, all operator paths will have the same number of $+i$ and $-i$ phase factors, cancelling out. Thus, all paths will have positive associated weights $W_{\mathcal{S}}$.
More generally, the sign problem is prevented as long as there are no closed loops generated by red or green transitions in the operator state graph. In the specific case of the mixed field QIM, shown in Fig. 3, adding a $h^{Y}Y_{i}$ term to the Hamiltonian will enable additional red and green transitions between states $X$ and $Z$ (via commutations with the $Y$ Pauli matrix). This will allow closed paths of the kind $X\rightarrow Y\rightarrow Z\rightarrow X$ with an $i^{3}$ phase to satisfy the fixed boundary conditions. A sufficient condition to prevent such closed loops (and consequently avoid the sign problem) is thus not to have all three Pauli matrices feature in the Hamiltonian expression. Beyond the mixed field QIM (which encompasses the Transverse Field Ising Model), the quantum XY model (in the absence of a $Z$ field term) is another notable quantum spin model that does not exhibit a sign problem. On the contrary, the quantum Heisenberg model, with its bond terms involving all three Pauli matrices, will suffer from the sign problem.
IV Numerical methods and observables
We now present a Monte Carlo sampling scheme for configurations in Eq. 8 for the mixed-field QIM. The moves are designed to traverse the configuration space by making local changes to the existing configuration. We describe them briefly here, and refer to Appendix A for technical details:
  1. 1.
    add-drop – allows switching between different $n$ sectors by adding or removing a pair of identical local Liouvillians located on the same sites at two consecutive time slices.
  2. 2.
    swap – swaps two commuting Liouvillians at different time slices, without changing the moment order $n$. This move is proposed only if there are no non-commuting Liouvillians at intermediate time slices.
  3. 3.
@@ IV Numerical methods and observables
    interact – here, two temporally consecutive $\mathcal{L}_{ZZ}$ Liouvillians that share a single spatial site “interact” to create a pair of temporally consecutive $\mathcal{L}_{Z}$ Liouvillians acting on next to nearest neighbour sites. This move allows accessing configurations with an odd number of $\mathcal{L}_{ZZ}$ and $\mathcal{L}_{Z}$ Liouvillians acting on a given site.
The weights $w_{n}$ are determined using a reweighting approach similar to the Wang-Landau algorithm [42, 43, 44]. An initial guess choice is iteratively tuned until reaching an approximately uniform distribution across the different $\mu_{2n}$ sectors. See Appendix A for implementation details.
@@
$r_{n}^{2}:=\frac{\mu_{2n+2}}{\mu_{2n}}=\frac{\left\langle\delta_{n^{\prime},n+1}\right\rangle_{n^{\prime}}}{\left\langle\delta_{n^{\prime},n}\right\rangle_{n^{\prime}}}\times\frac{w_{n}}{w_{n+1}}$ (10)
In the above equation, averaging is carried out over the probability distribution $P(n^{\prime})$ induced by $\mathcal{Z}$, where the Kronecker delta $\delta_{n^{\prime},n}$ counts Monte Carlo events satisfying $n^{\prime}=n$. The large frequency behavior of Eq. 1, predicts a linear growth of $r_{n}$ with $n$, i.e. $r_{n}\simeq 2\omega_{0}n$. In addition, the prefactor allows for a numerical estimation of the decay rate $\omega_{0}$.
@@
V Numerical Results
[…]
Figure 4: (a) $r_{n}$ vs $n$ for the $Z$ operator with $J=h^{Z}=h^{X}=1$ in 1D. The inset zooms in on the comparison of our MCMC calculations for the first 35 moments with brute-force exact numerics. (b) $\Delta r_{n}$ vs $n$ for 1D. This asymptotically falls to zero. Multiplying by $\log(2n)$ recovers a constant plateau. (c) Same as (a) for 2D. The inset compares the first 16 moments against exact computation. (d) $\Delta r_{n}$ vs $n$ for 2D.
Furthermore, we observe that the large $n$ growth of $r_{n}$ deviates from a linear trend. To pin down the source of this deviation and check whether it can be attributed to the aforementioned logarithmic correction in Fig. 4b, we plot $\Delta r_{n}\times\log(2n)$, where $\Delta r_{n}$ is a numerical derivative extracted using the five-point stencil method. This quantity is expected to reach a plateau at large $n$ with the logarithmic correction. Indeed, we find that only in the presence of the log multiplication, the desired plateau is obtained. By contrast, $\Delta r_{n}$ asymptotically decays to zero due to the $1/\log(2n)$ factor. We note that this is the first numerically exact confirmation for the predicted log correction associated with the restricted 1D dynamics, as previous brute-force numerics cannot resolve this subtle functional dependence convincingly due to the limited number of accessible moments [21, 37, 47].